Write the equation of the line that passes through point (–4, –16) with a slope of 3.
Found 2 solutions by chantal, checkley71:
Answer by chantal(1)

y=3x+-16
-16=3(-4)+b
-16=-12+b
+12 +12 (you add 12 to both sides of the equation
-4=b
so the equation of the line is y=3x-4

Answer by checkley71(8403)

Y=mX+b GIVEN A SLOPE OF 3 WE HAVE
Y=3X+b GIVEN X=-4 & Y=-16 WE HAVE
-16=3(-4)+b
-16=-12+b
b=-16+12
b=-4
THUS THE EQUATION FOR THIS LINE IS
Y=3X-4